# Nesterov Acceleration for Ensemble Kalman Inversion and Variants

### v1

@@ 1.1 Inverse problem
where $\eta$ is measurement noise, assumed to be Gaussian $\eta\sim\mathcal{N}(0,\Gamma)$. The inverse problem then is to estimate $u^{*}$ given $y$. We assume access to the model $\mathcal{G}$.
@@ 1.2 Ensemble Kalman inversion
We now turn to the estimation of $u_{\text{MLE}}$ using ensemble Kalman inversion [EKI, 1]. We note that while variants or extensions [2, 3, 4] of EKI can also be used to approximate the Bayesian inverse problem of estimating a posterior distribution $\mathbb{P}(u|y)$, here we restrict ourselves to finding a point estimate.
@@
We also note that (from [5]),
$C^{u\mathcal{G}}\approx C^{uu}\frac{d\mathcal{G}}{du}^{\top}.$ (4)
@@
Several variants of EKI exist. Unscented Kalman inversion [UKI, 9] chooses the ensemble using quadrature points at every iteration and has favorable performance for some problems. Ensemble transform Kalman inversion (ETKI), based on the ensemble transform Kalman filter [10], avoids building covariance matrices in the parameter and observation spaces, and has favorable computational complexity if $N\ll d,k$.
@@ 1.4 Motivation and related work
Nesterov acceleration for EKI was first suggested by [14]; however, detailed numerical tests were not conducted in that work, nor was Nesterov applied to EKI variants. Another method for accelerating convergence of EKI was considered in [15]. Momentum was considered for other particle-based optimization methods in [16], and for particle-based sampling methods in [17].
@@ 3.1 Example inverse problems
where the indices are cyclical and $x(t)=[x_{1}(t),\ldots,x_{D}(t)]^{\top}$. That is, we observe the state $x$ at $t_{K}$ after integrating from $x(t_{0})$. We use the fourth-order Runge–Kutta method with timestep $0.05$, and $t_{K}=0.4$. The dimension is set to $D=20$ and the forcing parameter to $F=8$, for which the Lorenz ’96 model is chaotic. To draw $u^{*}$, $D$ independent samples of $\mathcal{N}(0,1)$ are taken to be the state at $t_{0}-1000$, which are then run though the dynamics until $t_{0}$. The initial ensembles for $u$ are drawn from independent $\mathcal{N}(0,1)$ distributions in each dimension.
@@ 3.3 Results
Figure 1: Comparison of EKI convergence with and without acceleration, over the three inverse problems of Section 3.1, as measured by the logarithm of the cost function (2). The results were averaged over 50 trials for exponential sine and Lorenz ’96, and 10 trials for Darcy. Ribbons denote one standard error from the mean.
Figure 1 shows the results of applying Nesterov acceleration to EKI on the three inverse problems. The ensemble size is taken to be 10, 20, and 52, respectively, for the three inverse problems, weakly related to the increasing problem dimension (D shows that acceleration occurs regardless of ensemble size). Results are summarized over random trials where the data realization $y$ and the initial ensemble $\{u^{n}_{0}\}_{n=1}^{N}$ are both redrawn. In all three cases, Nesterov acceleration results in faster reduction of the cost function, as well as appearing to converge to better optima.
Figure 2: Comparison of EKTI and UKI convergence with and without acceleration, over the Exp Sin inverse problem of Section3.1, as measured by the logarithm of the cost function (2). Ribbons denote one standard error from the mean.
We now test Nesterov acceleration using UKI and ETKI (these algorithms are described in C) on the exponential sine problem. Figure 2 demonstrates that the performance is also improved with these methods. This suggests that Nesterov acceleration can be applied to a variety of EKI variants, such as those described in [9].
@@
In some of the cases, such as the UKI experiment in Figure 2, temporary “bumps” where the cost function increases with Nesterov-accelerated EKI can be observed. This may be because Nesterov-accelerated gradient descent, unlike gradient descent, is not strictly a descent method, in the sense that the objective function is not guaranteed to be monotonically non-increasing with iterations. This behavior is also observed when accelerating gradient descent [13].
@@ 5 Future work
Another avenue of future work is the extension of the current work to the ensemble Kalman sampler [EKS, 2]. Although there has been previous work on acceleration of continuous-time EKS [21, 17], the discrete-time versions of the algorithms are complex and depend in detail on the form of the EKS update. Our approach may offer simple implementations of such updates.
This work demonstrates that improvements of gradient descent algorithms can be transferred to EKI and its variants, by using the EKI’s continuous-time gradient flow limit as a bridge. This paves the way for other gradient-based optimizer advancements (e.g., Adam/Nadam/Adan [22, 23, 24]) to be transferred into the EKI framework.
@@ Appendix A Continuous-time limit for Nesterov-accelerated EKI
$(\bar{A}-\Delta t\bar{B})^{-1}=\bar{A}^{-1}+\sum_{k=0}^{\infty}(\Delta t)^{k}(\bar{A}^{-1}\bar{B})^{k}\bar{A}^{-1}=o(\Delta t).$
@@
where $c_{1}$, $c_{2}$, and $c_{3}$ are not functions of $\Delta t$. Then, defining $A=\Gamma+\Delta tC^{\mathcal{G}\mathcal{G}}(\{U^{(\cdot)}(t)\})$, we have for sufficiently small $\Delta t$ the Neumann series expansion
$(A+(\Delta t)^{3/2}c_{2}+o((\Delta t)^{3/2}))^{-1}$
$\qquad=\left(I+A^{-1}(\Delta t)^{3/2}c_{2}+o((\Delta t)^{3/2})\right)^{-1}A^{-1},$
$\qquad=(I-A^{-1}(\Delta t)^{3/2}c_{2}+o((\Delta t)^{3/2}))A^{-1},$
where we have used the fact that $A^{-1}$ is of order $\Delta t$ for sufficiently small $\Delta t$. Substituting back into (14) and expanding, we find the desired result (13).
Using the previous result that $A^{-1}$ is $o(\Delta t)$, we deduce

### v2

@@ 1.1 Inverse problem
where $\eta$ is measurement noise, assumed to be Gaussian $\eta\sim\mathcal{N}(0,\Gamma)$. The inverse problem is then to estimate $u^{*}$ given $y$. We assume access to the model $\mathcal{G}$. Note that $y$ can include multiple independent observations of the same quantities.
@@ 1.2 Ensemble Kalman inversion
We now turn to the estimation of $u_{\text{MLE}}$ using ensemble Kalman inversion [EKI, 1]. While we adopt the notation and terminology of [1], this methodology has a long lineage from randomized likelihood optimizers and ensemble smoothers [2, 3, 4, 5, 6] as used for state and parameter estimation in weather forecasting, and for oil reservoir history matching. Moreover, the idea of using an ensemble to approximate a gradient dates back to the introduction of the ensemble Kalman filter [7]. We note that while variants or extensions [8, 9, 10] of EKI can also be used to approximate the Bayesian inverse problem of estimating a posterior distribution $\mathbb{P}(u|y)$, here we restrict ourselves to finding a point estimate.
@@
We also note that under assumptions that the second derivative of $\mathcal{G}$ is small [11],
$C^{u\mathcal{G}}\approx C^{uu}\frac{d\mathcal{G}}{du}\Bigg{|}_{\mathbb{E}[u]}^{\top}.$ (4)
@@
We also note that EKI is often derived from the ensemble Kalman filter or from connection with the Bayesian solution of a linear inverse problem, wherein observations are perturbed for each ensemble member [1]. However, in this paper, since we use the formulation of EKI based on MLE, perturbed observations do not arise.
Several variants of EKI exist. Unscented Kalman inversion [UKI, 15] chooses the ensemble using quadrature points at every iteration and has favorable performance for some problems. Ensemble transform Kalman inversion (ETKI), based on the ensemble transform Kalman filter [16], avoids building covariance matrices in the parameter and observation spaces, and has favorable computational complexity if $N\ll d,k$.
@@ 1.4 Motivation and related work
Nesterov acceleration for EKI was first suggested by [20]; however, detailed numerical tests were not conducted in that work, nor was Nesterov applied to EKI variants. Another method for accelerating convergence of EKI was considered in [21]. A related implementation was also considered for the ensemble Kalman updates with a different objective [22]. There momentum was applied to the ensemble mean and covariance, and with constant momentum parameter. By contrast, we will showcase a simpler particle-wise update, and one with a momentum parameter that changes with iteration in a manner consistent with the gradient-based theory. Momentum has also been considered for particle-based sampling methods in [23].
@@ 3.1 Example inverse problems
where the indices are cyclical and $x(t)=[x_{1}(t),\ldots,x_{D}(t)]^{\top}$. We use the fourth-order Runge–Kutta method with timestep $0.05$, and $t_{K}=0.4$. The dimension is set to $D=20$ and the forcing parameter to $F=8$, for which the Lorenz ’96 model is chaotic. We draw the truth $u^{*}$ from the $D$-dimensional standard normal distribution. Taking $u^{*}$ to be the true state at time $t_{0}$, we run through the dynamics until time $t_{k}=t_{0}+1000$. The initial ensemble members for $u$ are each drawn from the $D$-dimensional standard normal distribution as well.
@@ 3.3 Results
Figure 1: Comparison of EKI, ETKI, and UKI convergence with and without acceleration, over the three inverse problems of Section 3.1, as measured by the logarithm of the cost function (2). The results were averaged over 50 trials for exponential sine and Lorenz ’96, and 10 trials for Darcy. Ribbons denote one standard error from the mean.
Figure 1 shows the results of applying Nesterov acceleration to the three algorithms EKI, ETKI, and UKI on the three inverse problems. For the first and second row, EKI and ETKI, the ensemble size is taken to be 10, 20, and 52, respectively, for the three inverse problems, weakly related to the increasing problem dimension (D shows that acceleration occurs regardless of ensemble size). The results are summarized over random trials where the data realization $y$ and the initial ensemble $\{u^{n}_{0}\}_{n=1}^{N}$ are both redrawn. In all three cases, and for all three algorithms, Nesterov acceleration results in faster reduction of the cost function, as well as appearing to converge to better optima. This suggests that Nesterov acceleration can successfully be applied to EKI variants, such as those described in [15].
@@
In some of the cases, such as the UKI experiment (bottom row of Figure 1), temporary “bumps” where the cost function increases with Nesterov-accelerated EKI can be observed. This may be because Nesterov-accelerated gradient descent, unlike gradient descent, is not strictly a descent method, in the sense that the objective function (2) is not guaranteed to be monotonically non-increasing with iterations. This behavior is also observed when accelerating gradient descent [19].
@@ 5 Future work
Another avenue of future work is the extension of the current work to the ensemble Kalman sampler [EKS, 8]. Although there has been previous work on acceleration of continuous-time EKS [27, 23], the discrete-time versions of the algorithms are complex and depend in detail on the form of the EKS update. Our approach may offer simple implementations of such updates, and more broadly, application to other empirically approximated non-Bayesian objective functions.
This work demonstrates that improvements of gradient descent algorithms can be transferred to EKI and its variants, by using the EKI’s continuous-time gradient flow limit as a bridge. This paves the way for other gradient-based optimizer advancements popular with the machine learning community to be transferred into the EKI framework. Examples of recently developed methods that could be implemented for EKI acceleration are the Information Theoretic Exact Method (ITEM) [28] or its cousin, triple momentum [29, 30] (corresponding to ITEM with constant momentum parameters); these methods have provable optimality. Additionally, there is the popular family of adaptive methods (namely, Adam, Nadam, Adan, Adamax, and Adagrad [31, 32, 33]) that lack in theoretical guarantees and continuum limits, but are used commonly in the machine learning community.
@@ Appendix A Continuous-time limit for Nesterov-accelerated EKI
$(A-\Delta tB)^{-1}=A^{-1}+\sum_{k=0}^{\infty}(\Delta t)^{k}(A^{-1}B)^{k}A^{-1}=o(\Delta t).$
@@
where $c_{1}$, $c_{2}$, and $c_{3}$ are not functions of $\Delta t$. Then, defining $D=\Gamma+\Delta tC^{\mathcal{G}\mathcal{G}}(\{U^{(\cdot)}(t)\})$, we have for sufficiently small $\Delta t$ the Neumann series expansion
$(D+(\Delta t)^{3/2}c_{2}+o((\Delta t)^{3/2}))^{-1}$
$\qquad=\left(I+D^{-1}(\Delta t)^{3/2}c_{2}+o((\Delta t)^{3/2})\right)^{-1}D^{-1},$
$\qquad=(I-D^{-1}(\Delta t)^{3/2}c_{2}+o((\Delta t)^{3/2}))D^{-1},$
where we have used the fact that $D^{-1}$ is of order $\Delta t$ for sufficiently small $\Delta t$. Substituting back into (14) and expanding, we find the desired result (13).
Using again the fact that $D^{-1}$ is $o(\Delta t)$ for $\Delta t$ sufficiently small, we deduce